The sum of two numbers is 70. The larger number is 4 more than twice the smaller number. What is the smaller number?
Found 2 solutions by solver91311, stanbon:
Answer by solver91311(24713)



Let represent the smaller number. Then must be the larger number. Then:



Solve for .

John

My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism


Answer by stanbon(75887)

The sum of two numbers is 70. The larger number is 4 more than twice the smaller number. What is the smaller number?
-----------------------
Equations:
x + y = 70
x = 2y+4
----
Substitute for "x" and solve for "y":
2y+4 + y = 70
3y = 66
y = 22
----
Solve for "x":
x = 2y+4
x = 2*22+4
x = 48
==============
Cheers,
Stan H.